### v1

@@ Abstract
Text-to-3D scene generation holds immense potential for the gaming, film, and architecture sectors. Despite significant progress, existing methods struggle with maintaining high quality, consistency, and editing flexibility. In this paper, we propose DreamScene, a 3D Gaussian-based novel text-to-3D scene generation framework, to tackle the aforementioned three challenges mainly via two strategies. First, DreamScene employs Formation Pattern Sampling (FPS), a multi-timestep sampling strategy guided by the formation patterns of 3D objects, to form fast, semantically rich, and high-quality representations. FPS uses 3D Gaussian filtering for optimization stability, and leverages reconstruction techniques to generate plausible textures. Second, DreamScene employs a progressive three-stage camera sampling strategy, specifically designed for both indoor and outdoor settings, to effectively ensure object-environment integration and scene-wide 3D consistency. Last, DreamScene enhances scene editing flexibility by integrating objects and environments, enabling targeted adjustments. Extensive experiments validate DreamScene’s superiority over current state-of-the-art techniques, heralding its wide-ranging potential for diverse applications. Code and demos will be released at https://dreamscene-project.github.io.
@@ 1 Introduction
We integrate optimized objects into the scene based on specific layouts to enhance scene generation, thereby preventing the production of duplicate or physically unrealistic artifacts, such as the multi-headed phenomena (generating the same content in all directions) in text-to-3D. Following scene generation, flexible scene editing can be achieved by individually adjusting objects and the environments (e.g., modifying positions and changing styles).
@@
  • Qualitative and quantitative experiments prove that DreamScene outperforms existing methods in text-driven 3D and 3D scene generation, unveiling substantial potential for numerous fields such as gaming, film, and architectural design.
@@ 2.1 Differentiable 3D Representation
Differentiable methods such as NeRF [21, 1, 23], SDF [26, 32], and 3D Gaussian Splatting [13] enable the representation, manipulation, and rendering of 3D objects and scenes. A significant advantage of these representations is their compatibility with optimization algorithms like gradient descent, facilitating the automatic tuning of 3D representation parameters to reduce loss, crucial for 3D reconstruction and generation tasks. Particularly, the recent approach [13] of modeling 3D scenes with differentiable 3D Gaussians—encompassing parameters such as position, covariance matrix, and opacity—has demonstrated enhanced rendering speed via splatting, achieving superior real-time rendering. Unlike implicit representations [21, 1, 23, 26, 32], 3D Gaussians provide a more explicit framework, simplifying the integration of multiple scenes. Consequently, this paper adopts 3D Gaussians for their clarity of explicit representation and ease of scene combination.
@@ 2.2 Text-to-3D Generation
Current text-to-3D tasks contain two main approaches: direct 3D representation generation [24, 12, 33] and 3D representation distilled from large 2D text-to-image models [28, 17, 2, 20, 10, 39, 16, 35, 15]. Direct generation methods, requiring annotated 3D datasets, facilitate rapid 3D representation generation. However, these methods often produce 3D structures of lower quality and demand significant GPU resources for training, commonly serving as initial steps for distillation techniques [16, 38]. For instance, Point-E [24] generates an image via a diffusion model [29, 30, 31] based on text, which is then transformed into the point cloud using a point cloud diffusion model. Shap-E [12] advances text-to-3D by mapping 3D assets to implicit function parameters with an encoder, followed by training a conditional diffusion model on these parameters.
Distilling 3D representation from 2D text-to-image models [28, 17, 2, 20, 10, 39, 16, 35, 15] has become the predominant strategy. DreamFusion [28], a trailblazer in this domain, introduced a Score Distillation Sampling (SDS) to ensure that images rendered from multiple perspectives conform to the distribution of 2D text-to-image models [29, 30, 31]. Following its lead, subsequent advancements [17, 2, 19, 20, 10, 39, 16, 35, 15, 24, 12] focused on refining 3D generation in terms of quality, speed, and diversity. For example, Magic3D [17] employs a low-resolution diffusion prior for initial rough model, then iteratively refines the textured 3D mesh model through interaction with an efficient differentiable renderer and a high-resolution latent diffusion model. DreamTime [10] accelerates the convergence of generation by monotonically non-increasing sampling of timesteps $t$ in 2D text-to-image model, while LucidDreamer [16] utilizes DDIM inversion [22, 6] to ensure 3D consistency of the object generation process. Inspired by these pioneering works, our method introduces a more efficient approach for generating high-quality and semantically rich 3D representation.
@@ 2.3 Text-to-3D Scene Generation Methods
Current text-to-3D scene generation methods, as illustrated in Fig. 1 face significant limitations. Methods [40, 9, 25] relying on inpainted images for scene completion can generate realistic visuals but suffer from limited 3D consistency of the generated scenes. Even minor disturbances can lead to the collapse of the scene’s visual integrity, indicating a lack of robustness. Moreover, such methods typically do not allow for the editing of objects within the scene and have unrealistic scene compositions, such as a living room cluttered with an excessive number of sofas. While some methods [3, 41], akin to ours, attempt to merge objects with environments for achieving a degree of consistency, they generally yield scenes of lower quality and can accommodate only a limited number of objects. Approaches [36, 43, 18] focusing solely on object integration fall short of generating comprehensive scenes. The constraints on quality, consistency, and editability have thus far hindered the broader application of text-to-3D technologies. Our method represents a leap forward by ensuring high-quality, scene-wide consistency and enabling flexible editing.
@@ 3 Preliminary
where $w(t)$ represents a weighting function that varies according to the timesteps $t$ and $y$ denotes the text embedding derived from the provided prompt. Classifier Score Distillation (CSD)  [39] is a variant of SDS, which, inspired by Classifier-Free Guidance (CFG) [8], divides the noise difference in SDS into a generation prior part $\epsilon_{\phi}(x_{t};y,t)-\epsilon$ and a classifier score part $\epsilon_{\phi}(x_{t};y,t)-\epsilon_{\phi}(x_{t};\emptyset,t)$. It posits that the classifier score is sufficiently effective for the text-to-3D part, it can be expressed as follows:
@@
3D Gaussian Splatting [13] is a cutting-edge technique in the field of 3D reconstruction. A 3D Gaussian is characterized by a full 3D covariance matrix $\Sigma$ defined in world space centered at point (mean) $\mu$:
@@
spherical harmonics (SH) coefficients, and opacity $\alpha$. By employing interlaced optimization and density control for these 3D Gaussians, particularly optimizing the anisotropic covariance, one can achieve precise representation of scenes. The implementation of a tile-based rendering strategy enables efficient anisotropic splatting, thereby accelerating the training process and achieving real-time rendering capability.
@@ 4 Method
The overarching concept of DreamScene involves dividing the scene into two components: objects and the environments. We first generate the objects and then place them within the environments to create the entire scene. This approach is designed to prevent the generation of objects in the environments and to mitigate the multi-head problem in 3D generation (outdoor environments can be generated separately). Specifically, we utilize a Large Language Model (LLM) [42] to decompose scene descriptions, obtaining descriptions for each object $y_{i}(i=1,..,N)$ and the environment $y_{e}$, and rapidly create high-quality objects using Formation Pattern Sampling. Then, we initialize a cuboid 3D Gaussians to simulate walls, floors, and ceilings for bounded indoor scenes and a hemispherical 3D Gaussians to simulate the ground and faraway surroundings for outdoor unbounded scenes. Through scene layout diagrams, we designate the positions of objects within the scene and use affine transformations (scale $s$, translation $t$, rotation $r$) to place objects in the scene coordinate system.
@@
where $x$ denotes the collection of coordinates for 3D Gaussians under the object’s coordinate. Finally, we employ a camera sampling strategy for multi-stage optimization of the environments, achieving scene generation with high 3D consistency.
@@ 4.1 Formation Pattern Sampling
We refined and broadened the concept of using monotonically non-increasing sampling of timesteps $t$ from DreamTime [10]. We find that the development of high-quality, semantically rich 3D representations benefits from integrating information across multiple timesteps of a pre-trained 2D text-to-image diffusion model at each iteration. Specifically, during the initial and middle phases of optimization, which focus on the primary shape formation of the 3D representation, we establish a decremental time window $T_{end}$ that linearly decreases with iterations. $T_{end}$ is segmented into $m$ intervals, within each of which $t$ is randomly sampled and gradients are aggregated. This technique swiftly yields semantically rich 3D representations but may accumulate superfluous 3D Gaussians. To address this, we apply 3D Gaussian filtering to preserve only the essential surface Gaussians. In the subsequent stages, focusing on surface texture refinement, we restrict $t$ to the range of $0\sim 200$ and employ 3D reconstruction methods [13] to expedite the generation of plausible textures.
We leverage pseudo-Ground-Truth (pseudo-GT) images, derived from a single denoising step in LucidDreamer [16], to encapsulate the diverse information provided by different timesteps $t(0\sim 1000)$ of the 2D text-to-image diffusion model. By adding $t$ timesteps of noise to images $x_{0}$ to obtain $x_{t}$, we estimate the pseudo-GT $\hat{x}_{0}^{t}$ by:
$\hat{x}_{0}^{t}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{t},t,y)}{\sqrt{\bar{\alpha}_{t}}}.$ (7)
@@
As shown in Fig. 3(a), we discovered when timestep $t$ is small, the 2D diffusion model provides the information of surface textures that are more detailed and realistic, and shape consistent with the current 3D representation, but lacks in providing rich semantic information of prompt $y$. When $t$ is large, the 2D diffusion model offers abundant semantic information, but its shape may not align with the current 3D representation (e.g., the orientation of the man, the color of the chair, the direction of the cooker between $t$ values of 600 to 800). Continuously using smaller $t$ tends to make the generated object lean towards simple semantic information, resulting in a significant gap from the prompt $y$. Conversely, continuously using larger $t$ leads to a lack of uniformity in the optimization direction of the 3D representation with each iteration, causing the shape to collapse.
Therefore, we propose to hybridize information from different 2D diffusion timesteps in each iteration, to ensure a certain degree of shape constraint while enriching the semantic information. For instance, in Fig. 3(a), at the $300$-th iteration phase of the man, we need to restrict the shape using the information of timesteps in $200\sim 400$, get richer semantic information using the information of timesteps in $400\sim 600$ and $600\sim 800$. However, at the $1000$-th iteration of the cooker, we find that the current 3D representation already possesses ample semantic information, and the information provided by larger $t$ might hinder the current optimization direction. Therefore, similar to DreamTime [10], our $i$-th sampled t can be expressed as:
@@
where $T_{end}$ denotes a linear decreasingly time window, $iter$ represents the current iterations of the optimzation process, and $m$ denotes the number of intervals. The multi-timestep sampling (MTS) combined with the CSD (Classifier Score Distillation) [39] method can be expressed as:
$\nabla_{\theta}\mathcal{L}_{\text{MTS}}(\theta)=\mathbb{E}_{t,\epsilon,c}\left[\sum\limits_{i=1}^{m}w(t)(\epsilon_{\phi}(x_{t_{i}};y,t_{i})-\epsilon_{\phi}(x_{t_{i}};\emptyset,t_{i}))\frac{\partial g(\theta,c)}{\partial\theta}\right].$ (9)
@@
Excessive Gaussians can hinder the optimization process. Unlike current methods [4, 14] that filter reconstructed 3D Gaussians using ground truth images, our approach necessitates filtering during the optimization process. In the context of rendering, 3D Gaussians positioned nearer to the rendering plane inherently have a more significant impact on the rendered image. To quantify this impact, we employ a score function tailored to capture such contributions. For 3D Gaussians along the rendering ray $r_{j}$, their contributions are calculated using the inverse square of their distance to the rendering plane, adjusted by the 3D Gaussians volume. This method prioritizes 3D Gaussians that are both closer to the rendering plane and larger in volume as shown in Fig. 8. By ranking various viewpoints according to their scores, we can efficiently eliminate 3D Gaussians falling below a designated threshold. The scores are calculated as follows:
@@
where $H$ represents the height of the rendered image, $W$ denotes the width, $M$ indicates the number of rendered images, $V(i)$ represents the volume of the $i$-th 3D Gaussian (calculated using the covariance matrix $\Sigma$), $maxV(r_{j})$ represents the maximum volume of the 3D Gaussians on $r_{j}$, and $D(r_{j},i)$ denotes the distance of the $i$-th 3D Gaussian from the rendering plane along the $r_{j}$. It’s important to note that this procedure mirrors the rendering process instead of performing the actual rendering.
@@
In the later stages of the optimization, we can expedite the generation of plausible surface textures for the 3D representation using 3D reconstruction methods [13]. During the previous optimization process, we find that when sampling very small timesteps $t$($0\sim 200$), the image $\hat{x}_{0}^{t}$ predicted by Eq. 7 has no difference in 3D shape compared to the image $x_{0}$ before noise addition, but offers more detailed and plausible textures. Therefore, under the premise of shape consistency, we can directly generate a new 3D representation through 3D reconstruction [13]. As shown in Fig. 3(c), after obtaining a coarse texture but highly consistent 3D representation in the previous step, we render $K$ images $x_{i},i=1,...,K$ from various camera poses $c_{i}$ around the 3D representation. By adding $t$ timesteps of noise to the images to get $x_{it}$ by Eq. 1, we estimate the images $\hat{x}_{i0}^{t}$ with plausible textures using Eq. 7, and then reconstruct them on the coarse representation through 3D reconstruction methods [13] by minimize the following reconstruction loss:
@@
This process generates a representation with detailed and plausible textures using only 15 seconds.

### v2

@@ Abstract
Text-to-3D scene generation holds immense potential for the gaming, film, and architecture sectors. Despite significant progress, existing methods struggle with maintaining high quality, consistency, and editing flexibility. In this paper, we propose DreamScene, a 3D Gaussian-based novel text-to-3D scene generation framework, to tackle the aforementioned three challenges mainly via two strategies. First, DreamScene employs Formation Pattern Sampling (FPS), a multi-timestep sampling strategy guided by the formation patterns of 3D objects, to form fast, semantically rich, and high-quality representations. FPS uses 3D Gaussian filtering for optimization stability, and leverages reconstruction techniques to generate plausible textures. Second, DreamScene employs a progressive three-stage camera sampling strategy, specifically designed for both indoor and outdoor settings, to effectively ensure object and environment integration and scene-wide 3D consistency. Last, DreamScene enhances scene editing flexibility by integrating objects and environments, enabling targeted adjustments. Extensive experiments validate DreamScene’s superiority over current state-of-the-art techniques, heralding its wide-ranging potential for diverse applications. Code and demos are released at https://dreamscene-project.github.io.
@@ 1 Introduction
We integrate optimized objects into the scene based on specific layouts to enhance scene generation, thereby preventing the production of duplicate or physically unrealistic artifacts, such as the multi-headed phenomena (generating the same content in all directions) in text-to-3D. Following scene generation, flexible scene editing can be achieved by individually adjusting objects and environments (e.g., modifying positions and changing styles).
@@
  • Qualitative and quantitative experiments prove that DreamScene outperforms existing methods in text-driven 3D object and scene generation, unveiling substantial potential for numerous fields such as gaming and film.
@@ 2.1 Differentiable 3D Representation
Differentiable methods such as NeRF [22, 1], SDF [27, 33], and 3D Gaussian Splatting [14] enable the representation, manipulation, and rendering of 3D objects and scenes. These representations are compatible with optimization algorithms like gradient descent, facilitating the automatic tuning of 3D representation parameters to reduce loss. Particularly, the recent approach [14] of modeling 3D scenes with differentiable 3D Gaussians has achieved superior real-time rendering via splatting. Unlike implicit representations [22, 1, 24], 3D Gaussians provide a more explicit framework, simplifying the integration of multiple scenes. Hence, we adopt 3D Gaussians for their clarity of explicit representation and ease of scene combination.
@@ 2.2 Text-to-3D Generation
Current text-to-3D tasks mainly generate 3D representation directly [25, 13, 34] or distilled from large 2D text-to-image models [29, 18, 3]. Direct methods require annotated 3D datasets for rapid generation, often suffering from lower quality and high GPU demands, thus commonly serving as initial steps for distillation techniques [17, 39]. For instance, Point-E [25] generates an image via a diffusion model based on text, which is then transformed into the point cloud. Shap-E [13] maps 3D assets to implicit function parameters with an encoder, then trains a conditional diffusion model on these parameters.
Distilling 3D representation from 2D text-to-image models [29, 18, 3, 21, 11, 40] has become the predominant strategy. DreamFusion [29], a trailblazer in this domain, introduced Score Distillation Sampling (SDS) to ensure that images rendered from multiple perspectives conform to the distribution of 2D text-to-image models [30, 31, 32]. Following its lead, subsequent advancements [18, 3, 20, 21, 11, 40, 17, 36, 16, 25, 13] refined 3D generation in terms of quality, speed, and diversity. For example, DreamTime [11] accelerates the convergence of generation by monotonically non-increasing sampling of timesteps $t$ in 2D text-to-image model, while LucidDreamer [17] utilizes DDIM inversion [23, 7] to ensure 3D consistency of the object generation process. Inspired by these pioneering works, our method introduces a more efficient approach for generating high-quality and semantically rich 3D representation.
@@ 2.3 Text-to-3D Scene Generation Methods
Current text-to-3D scene generation methods, as illustrated in Fig. 1 face significant limitations. [41, 10, 26] relying on inpainted images for scene completion can generate realistic visuals but suffer from limited 3D consistency. Even minor disturbances can lead to the collapse of the scene’s visual integrity, indicating a lack of robustness. Moreover, such methods typically do not allow for the editing of objects within the scene and have unrealistic scene compositions, such as a living room cluttered with an excessive number of sofas. While some methods [4, 42], akin to ours, attempt to merge objects with environments to ensure consistency, they generally yield scenes of lower quality and can accommodate only a limited number of objects. Approaches [37, 44, 19] focusing solely on object integration fall short of generating comprehensive scenes. The constraints on quality, consistency, and editability have thus far hindered the broader application of text-to-3D technologies. Our method represents a leap forward by ensuring high-quality, scene-wide consistency and enabling flexible editing.
@@ 3 Preliminary
where $w(t)$ represents a weighting function that varies according to the timesteps $t$ and $y$ denotes the text embedding derived from the provided prompt. Classifier Score Distillation [40] is a variant of SDS inspired by Classifier-Free Guidance (CFG) [9], divides the noise difference in SDS into generation prior $\epsilon_{\phi}(x_{t};y,t)-\epsilon$ and classifier score $\epsilon_{\phi}(x_{t};y,t)-\epsilon_{\phi}(x_{t};\emptyset,t)$. It posits that the classifier score is sufficiently effective for text-to-3D, expressed as follows:
@@
3D Gaussian Splatting [14, 2] is a cutting-edge technique in the field of 3D reconstruction. A 3D Gaussian is characterized by a full 3D covariance matrix $\Sigma$ defined in world space centered at point (mean) $\mu$:
@@
spherical harmonics (SH) coefficients, and opacity $\alpha$. By employing interlaced optimization and density control for these 3D Gaussians, particularly optimizing the anisotropic covariance, one can achieve precise scene representation. The tile-based rendering strategy enables efficient anisotropic splatting, thereby accelerating the training process and achieving real-time rendering capability.
@@ 4 Method
DreamScene divides the scene into objects and environments. It first generates the objects and then places them within the environments to create the entire scene. This approach is designed to prevent the generation of objects in the environments and to mitigate the multi-head problem in 3D generation (outdoor environments can be generated separately). Specifically, we employ prompt engineering to develop a large language model (LLM) [43] agent that transforms scene text into detailed object descriptions $y_{i}(i=1,..,N)$ and environment description $y_{e}$, and rapidly create high-quality objects using Formation Pattern Sampling, which consists of multi-timestep sampling, 3D Gaussian filtering and reconstructive generation. Then, we initialize cuboid 3D Gaussians to simulate walls, floors, and ceilings for bounded indoor scenes and hemispherical 3D Gaussians to simulate the ground and faraway surroundings for outdoor unbounded scenes. Through scene layout diagrams, we designate the positions of objects within the scene and use affine transformations (scale $s$, translation $t$, rotation $r$) to place objects in the scene coordinate system.
@@
where $x$ denotes the collection of coordinates for 3D Gaussians. Finally, we employ a camera sampling strategy for multi-stage optimization of the environments, achieving scene generation with high 3D consistency.
@@ 4.1 Formation Pattern Sampling
We have refined and broadened the concept of using monotonically non-increasing sampling of timesteps $t$ from DreamTime [11]. We find that the development of high-quality, semantically rich 3D representations benefits from integrating information across sampling multiple timesteps of a pre-trained 2D text-to-image diffusion model at each iteration. This finding contrasts with all other methods using SDS [29], which typically utilize information from a single timestep at each iteration. Specifically, during the initial and middle phases of optimization, which focus on the primary shape formation, we establish a decremental time window $T_{end}$ that linearly decreases with iterations. $T_{end}$ is segmented into $m$ intervals, within each of which $t$ is randomly sampled and gradients are aggregated. This technique swiftly yields semantically rich 3D representations but may accumulate superfluous 3D Gaussians. To address this, we apply 3D Gaussian filtering to sample only the essential surface Gaussians. In the subsequent stages, focusing on surface texture refinement, we sample $t$ within the range of $0\sim 200$ and employ 3D reconstruction methods [14] to expedite the generation of plausible textures. In this process, we follow the pattern of 3D model formation to sample different timesteps $t$ at different iterations and 3D Gaussians on the model surface, hence we refer to this process as Formation Pattern Sampling. We leverage pseudo-Ground-Truth (pseudo-GT) images, derived from a single denoising step in LucidDreamer [17], to encapsulate the diverse information provided by different timesteps $t(0\sim 1000)$ of the 2D text-to-image diffusion model. By adding $t$ timesteps of noise to images $x_{0}$ to obtain $x_{t}$, we estimate the pseudo-GT $\hat{x}_{0}^{t}$ by:
$\hat{x}_{0}^{t}=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{t};y,t)}{\sqrt{\bar{\alpha}_{t}}}.$ (7)
@@
As shown in Fig. 3(a), we discovered when timestep $t$ is small, the 2D diffusion model provides more detailed and realistic surface textures, of which the shape is consistent with the current 3D representation, but lacks in providing rich semantic information of prompt $y$. When $t$ is large, the 2D diffusion model offers abundant semantic information, but its shape may not align with the current 3D representation (e.g., the orientation of the man, the color of the chair, the direction of the cooker between $t$ values of 600 to 800). Therefore, we propose to hybridize information from different 2D diffusion timesteps in each iteration, to ensure a certain degree of shape constraint while enriching the semantic information. For instance, in Fig. 3(a), at the $300$-th iteration phase of the man, we need to restrict the shape using the information of timesteps in $200\sim 400$, get richer semantic information using the information of timesteps in $400\sim 600$ and $600\sim 800$. However, at the $1000$-th iteration of the cooker, we find that the current 3D representation already possesses ample semantic information, and the information provided by larger $t$ might hinder the current optimization direction. Our $i$-th sampled t can be expressed as:
@@
where $T_{end}$ denotes a linear decreasingly time window similar to DreamTime, $iter$ represents the current iteration, and $m$ denotes the number of intervals. We adopt DDIM Inversion to sample from $t_{1}$ to $t_{m}$ to ensure content consistency [17].
Therefore, the multi-timestep sampling (MTS) combined with the CSD (Classifier Score Distillation) [40] method can be expressed as:
$\nabla_{\theta}\mathcal{L}_{\text{MTS}}(\theta)=\mathbb{E}_{t,\epsilon,c}\left[\sum\limits_{i=1}^{m}w(t_{i})(\epsilon_{\phi}(x_{t_{i}};y,t_{i})-\epsilon_{\phi}(x_{t_{i}};\emptyset,t_{i}))\frac{\partial g(\theta,c)}{\partial\theta}\right].$ (9)
@@
Excessive Gaussians can hinder the optimization process. Unlike current methods [5, 15] that filter reconstructed 3D Gaussians using ground truth images, our approach necessitates filtering during optimization. In terms of rendering, the 3D Gaussians closer to the rendering plane inherently have a more significant impact, for which we employ a score function tailored to quantify such contributions. For 3D Gaussians along the rendering ray $r_{j}$, their contributions are calculated using the inverse square of their distance to the rendering plane, adjusted by the 3D Gaussians volume. This method prioritizes 3D Gaussians that are both closer to the rendering plane and larger in volume as shown in Fig. 8. By ranking various viewpoints according to scores, we can efficiently eliminate 3D Gaussians falling below a designated threshold.
@@
where $H$ and $W$ denote the height and width of the rendered image, $M$ indicates the number of rendered images, $V(i)$ represents the volume of the $i$-th 3D Gaussian (calculated using the covariance matrix), $maxV(r_{j})$ represents the maximum volume of the 3D Gaussians on $r_{j}$, and $D(r_{j},i)$ is the distance of the $i$-th 3D Gaussian from the rendering plane along the $r_{j}$. Note that this procedure mirrors the rendering process instead of performing the actual rendering.
@@
Next, we can expedite the generation of plausible surface textures using 3D reconstruction methods [14]. During the previous process, we find that when sampling very small timesteps $t$($0\sim 200$), the image $\hat{x}_{0}^{t}$ predicted by Eq. 7 has no difference in 3D shape compared to the image $x_{0}$ before noise addition, but offers more detailed and plausible textures. Therefore, under the premise of shape consistency, we can directly generate a new 3D representation through 3D reconstruction [14]. As shown in Fig. 3(c), after obtaining a coarse texture and highly consistent 3D representation, we render $K$ images $x_{i},i=1,...,K$ from various camera poses $c_{i}$ around the 3D representation. By adding $t$ timesteps of noise to the images to get $x_{it}$ by Eq. 1, we estimate the images $\hat{x}_{i0}^{t}$ with plausible textures using Eq. 7, and then reconstruct them on the coarse representation through [14] by minimizing the following reconstruction loss:
@@
This process generates a representation with detailed and plausible textures within 15 seconds.